# Jasper-Token-Compression-600M Technical Report

### v1

@@ 2.1 Knowledge Distillation
Because the embedding dimensions of Qwen3-Embedding-8B and QZhou-Embedding are 4096 and 3584, respectively, we apply special processing to facilitate subsequent distillation. Given that Qwen3-Embedding-8B adopts Matryoshka Representation Learning (MRL) 10.5555/3600270.3602462 during training, allowing custom output dimensions for the final embedding, we directly take the first 1024 dimensions of its embedding, denoted as $E_{qwen}$. For QZhou-Embedding which does not support MRL natively, we take the first 3072 dimensions of its 3584-dimensional embedding, split them into three contiguous 1024-dimensional segments, and sum the segments element-wise to obtain a 1024-dimensional vector, denoted as $E_{qzhou}$. The final 2048-dimensional teacher embedding is then constructed as: $E_{t}=\text{Norm}(\text{Norm}(E_{\text{qwen}})\,|\,\text{Norm}(E_{\text{qzhou}}))$ , where “||” denotes concatenation along the feature dimension, and “Norm” denotes L2 normalization. The original Qwen3-Embedding-0.6B produces a 1024-dimensional embedding based on last-token pooling. We modify this to mean pooling and introduce a randomly initialized linear layer to map the 1024-dimensional vector to a 2048-dimensional student embedding. After L2 normalization, the student embedding is denoted as $E_{s}$. During this stage, we adopt the classical cosine loss, formulated as:
@@ 2.2 Distillation with Token Compression
We perform distillation training using the same data as in the first stage. During this stage, all parameters of the student model are trained. It is important to note that the newly introduced token compression module only has the Qwen3MLP as trainable, while the AdaptiveAvgPool1d is training-free. We train for 2 epochs using the Adam optimizer with a learning rate of 7e-5, with a length threshold of 80 and a compression ratio of 0.33. All other parameters are kept the same as in the first stage.
@@ 2.4 Contrastive Learning for Retrieval
After the three-stage distillation procedure described above, we observed that the student model achieves performance comparable to the teacher models on most MTEB tasks. However, its retrieval performance still lag behind those of the teacher models. To address this gap, we introduce the fourth contrastive learning-based training stage, specifically designed to enhance retrieval capacities.
We adopt an improved contrastive objective built upon the InfoNCE framework oord2019representationlearningcontrastivepredictive. Given a batch of $N$ training instances, the contrastive loss is defined as:
@@
To further guide the student model during contrastive learning, we additionally incorporate a soft distillation loss:
@@
where $S^{(s)}$ and $S^{(t)}$ denote the similarity score vectors computed from the student and teacher embeddings (each of length $N(1+K)$), respectively, $\mathrm{sm}(\cdot)$ denotes the softmax function, and $\alpha$ is a temperature scaling factor used to adjust similarity differences. The final loss function for Stage 4 is therefore:
$\mathcal{L}_{\text{s4}}=L_{\text{cl}}+16\cdot L_{\text{soft}}.$ (7)
@@ 3.2 Ablation Study
As discussed in Section 2.4, we introduce a fourth training stage based on contrastive learning to specifically enhance the student model’s retrieval capability. As shown in Table 3, the Stage 3 student model already matches the teacher model (Qwen3-Embedding-8B) on tasks such as Classification and Clustering. However, it still lags behind the teacher by nearly 4 points on asymmetric Retrieval tasks (65.53 vs. 69.44). After applying the retrieval-oriented contrastive learning in Stage 4, we observe a modest improvement on the Retrieval task (from 65.33 to 66.19), while performance on other tasks fluctuates only slightly. This result indicates that targeted contrastive learning is indeed effective for improving retrieval capability.

### v2

@@ 2.1 Knowledge Distillation
Given that the embedding dimensions of Qwen3-Embedding-8B and QZhou-Embedding are 4096 and 3584, respectively, we apply special processing to facilitate subsequent distillation. Given that Qwen3-Embedding-8B adopts Matryoshka Representation Learning (MRL) 10.5555/3600270.3602462 during training, allowing custom output dimensions for the final embedding, we directly take the first 1024 dimensions of its embedding, denoted as $E_{qwen}$. For QZhou-Embedding which does not support MRL natively, we take the first 3072 dimensions of its 3584-dimensional embedding, split them into three contiguous 1024-dimensional segments, and sum the segments element-wise to obtain a 1024-dimensional vector, denoted as $E_{qzhou}$. The final 2048-dimensional teacher embedding is then constructed as: $E_{t}=\text{Norm}(\text{Norm}(E_{\text{qwen}})\,|\,\text{Norm}(E_{\text{qzhou}}))$ , where “||” denotes concatenation along the feature dimension, and “Norm” denotes L2 normalization. The original Qwen3-Embedding-0.6B produces a 1024-dimensional embedding based on last-token pooling. We modify this to mean pooling and introduce a randomly initialized linear layer to map the 1024-dimensional vector to a 2048-dimensional student embedding. After L2 normalization, the student embedding is denoted as $E_{s}$. During this stage, we adopt the classical cosine loss, formulated as:
@@ 2.2 Distillation with Token Compression
We perform distillation training using the same data and loss function as in the first stage. During this stage, all parameters of the student model are trained. It is important to note that the newly introduced token compression module only has the Qwen3MLP as trainable, while the AdaptiveAvgPool1d is training-free. We train for 2 epochs using the Adam optimizer with a learning rate of 7e-5, with a length threshold of 80 and a compression ratio of 0.33. All other parameters are kept the same as in the first stage.
@@ 2.4 Contrastive Learning for Retrieval
After the three-stage distillation procedure described above, we observed that the student model achieves performance comparable to the teacher models on most MTEB tasks. However, its retrieval performance still lag behind those of the teacher models.
To address this gap, we introduce the fourth contrastive learning-based training stage, specifically designed to enhance retrieval ability. We adopt an improved contrastive objective built upon the InfoNCE framework oord2019representationlearningcontrastivepredictive. Given a batch of $N$ training instances, the contrastive loss is defined as:
@@
To effectively guide the student model during contrastive learning, we further incorporate a soft distillation loss:
@@
where $S^{(s)}$ and $S^{(t)}$ denote the similarity score vectors computed from the student and teacher embeddings (each of length $N(1+K)$), respectively, $\mathrm{sm}(\cdot)$ denotes the softmax function, and $\alpha$ is a temperature scaling factor used to adjust similarity differences. Additionally, we employ the $\mathcal{L}_{\text{cosine}}$ previously used in the three training stages as a supplementary regularization term. The final loss function for Stage 4 is therefore:
$\mathcal{L}_{\text{s4}}=\mathcal{L}_{\text{cl}}+16\cdot\mathcal{L}_{\text{soft}}+10\cdot\mathcal{L}_{\text{cosine}}.$ (7)
@@ 3.2 Ablation Study
As discussed in Section 2.4, we introduce the fourth training stage based on contrastive learning to specifically enhance the student model’s retrieval capability. As shown in Table 3, the Stage 3 student model already matches the teacher model (Qwen3-Embedding-8B) on tasks such as Classification and Clustering. However, it still lags behind the teacher by nearly 4 points on asymmetric Retrieval tasks (65.53 vs. 69.44). After applying the retrieval-oriented contrastive learning in Stage 4, we observe a modest improvement on the Retrieval task (from 65.33 to 66.19), while performance on other tasks fluctuates only slightly. This result indicates that targeted contrastive learning is indeed effective for improving retrieval capability.